There is problem with 3over 5y=6 and I don't know the process to solce
Answer by MathDazed(34)

I am thinking you are writing 3/5y = 6
What you do is multiply both sides by 5y
(5y)*3/5y = 6*5y
3 = 30y
Then divide both sides by 30
3/30 = 30y/30
3/30 = y
Simplify
1/10 = y